# Physics Informed Generative Machine Learning for Accelerated Quantum-centric Supercomputing

### v1

@@ III.2 Generative Machine Learning Based Configuration Subspace Expansion
where, is the partition function. Since explicit calculation of the partition function becomes computationally intractable, RBMs are usually trained on a set of vectors distributed according to some unknown probability distribution using the contrastive divergence (CD) method which approximates the gradient of the log likelihood to learn the underlying distribution. RBM alternatively updates hidden and visible layers using conditional activation probabilities:
where, is the sigmoid function. By maximizing the log likelihood of the data , RBM uses stochastic approximation to update its parameters which is equivalent to minimizing the Kullback-Leibler divergence between the data distribution and model distribution. After training, a new set of vectors are generated according to the learnt distribution using Gibbs sampling[53] technique. For a detailed description of the working principle of RBM, the readers are referred to Refs.[54, 55, 56, 57].
@@
Initial State Preparation and Sampling from Quantum Hardware: State preparation is done on a quantum hardware with an ansatz that has a substantial overlap with the ground state. Usually a PQC corresponding to a unitary is prepared that acts on HF state such that the prepared state . This is followed by repeated measurements performed with a specific number of shots to obtain samples from the hardware that produce a measurement outcome with a probability distribution where each bitstring is a computational basis state.
@@
Micro Cycle: Each macro-cycle contains a set of micro-cycle iterations which is used to train RBM with the prepared distribution in the previous step. Trained RBM is used to generate new configurations using Gibbs sampling method. Toward this, a set of spin-particle conserved configurations, whose number is fixed at % of , are randomly sampled from the symmetry space and given to the Gibbs sampler which generate same number of spin-particle symmetry conserved configurations.
@@ IV Computational Details
All the numerical results in this section are obtained by in-house python based codes. MBPT calculations are done using two-body integrals and MP2 amplitudes obtained from PySCF[58] driver in Qiskit[59]. Three strongly correlated molecular systems , and in 6-31G basis with frozen core approximation involving 24, 32 and 40 spin orbitals correlating 8, 10 and 10 electrons respectively are studied in this work. The FCI, CCSD, CI with singles and doubles (CISD), MP2 and HF calculations were performed using PySCF with the frozen core approximation. Since FCI calculations are not feasible for , heat-bath configuration interaction (HCI)[33] calculations were done using PySCF to get reference energy values for accuracy comparison. The entire workflow involves multiple independent computational components orchestrated together as discussed below:
Quantum state preparation and hardware sample generation: State preparation on a quantum hardware is done using the local unitary coupled jastrow (LUCJ) ansatz [49] (Eq. (6)) with CCSD parameters used for jastrow decomposition. No further optimization of the LUCJ parameters is performed. For 24 qubits experiment 250000 shots were used while both 32 and 40 qubit experiments were performed using shots. The circuits were created using qiskit-addon-sqd, which uses the ffsim[60] package to generate the LUCJ circuits. For each circuit, we start with some generic linear layout and invoke the VF2PostLayout transpiler pass to find an isomorphic zig-zag layout (which is required for LUCJ circuits) that conforms to the lowest noise profile as obtained from the backend noise information. The LUCJ circuit is placed on this layout using the qiskit transpiler. We executed all the circuits on ibm_kingston, which is a 156 qubits IBM Heron R2 processor. The layouts used for the circuits corresponding to each of the molecules are shown in Fig. 1.
@@
To obtain the physics-informed support to ground state via perturbative measures, initially available molecular integrals (in molecular orbital basis) are filtered from the original sparse integral tensor using a threshold . All surviving integral values, together with their associated index combinations, are collected and arranged into an array to construct following the integral index pattern (). The computational cost of triples selection step is governed by the most expensive term which can be performed efficiently using the symbolic index operation technique discussed in Sec.III. For this efficient generation we combine all unique outer indices and having only one index common between them (see Appendix B) with the constraints and respectively, stemming from pruned and . Such an operation does not require explicit summations over contractible indices, and just one common index suffices the inclusion of the corresponding outer (non-contractible) indices. At this stage, it is important to note that the hierarchical process of configuration selection through perturbation theory automatically ensures that the lower order terms, double excitations () from the HF reference, are already pruned which implies . This provides the flexibility to avoid the prohibitive cost of explicit tensor contraction, saving huge number of operations to be performed. A comparison of the number of operations required for the pruned elements in the numerical worst case scenario for triples and quadruples selection are shown in Fig.2 against the formal asymptotic scaling of , CCSD and CCSDT. Fig. 2 shows the numerical limiting case of symbolic tensor manipulation with pruned elements for triples selection requires almost two orders-of-magnitude less number of operations compared to the asymptotic cost of explicit computations. The actual implementational cost for the same using the symbolic technique requires much less number of operations than this in all practical cases. A similar approach can be followed for constructing perturbative quadruples as well by breaking the associated contraction into intermediates and performing the index-based symbolic calculations. The algebraic terms calculated for selection of quadruply excited configurations are given in Appendix A. The corresponding numerical limiting case scales almost similar to CCSD as demonstrated in Fig.2.
@@
While this way of selection of important high-rank excited configurations through their lowest order perturbative measures is not theoretically exact from a conventional many-body viewpoint, we argue that an absolute accuracy (of choosing the important configurations) is not also warranted for; rather, this measure gives us a strong support to the hardware generated configuration samples. Dominant determinants that are missed out (due to noise or perturbatively) would eventually get picked up through generative ML. Also, in a similar spirit, any determinant which are spuriously picked up would get dropped (blacklisted) during the iterative fermionic state reconstruction as will be shown in numerical demonstrations.
@@
Generative ML training and fermionic state reconstruction: In this study, we have used RBM with Contrastive Divergence (CD) algorithm for training. The dimensions of visible and hidden layer are kept the same in this study which is equal to the number of spin orbitals, the learning rate is fixed at 0.001. The training distribution provided to the RBM is constructed by inflating the current configuration space obtained in a given macro-iteration, with each configuration replicated in proportion to its associated CI coefficient (excluding the HF configuration, as discussed in Sec. III.2). Within each RBM micro-cycle, three training iterations are performed using the CD algorithm to update the model parameters. During these updates, CD incorporates an intrinsic Gibbs-sampling step, for which we employ 20 Gibbs iterations. During generation, only one Gibbs sampling step is used that takes a fixed number of randomly selected configurations from the correct spin-particle sector of the Hilbert space and generates the same number of configurations using the learned parameters. The number of samples used for Gibbs sampling and subsequent generation at each macro-cycle is fixed a priori by configurations of the symmetry space where the number is specified for the numerical studies in the result section (Sec. V). Such a training and generation process avoids over-training and retain a low degree of stochasticity in the model for better generation of new configurations as discussed further in Sec.V.
@@ V Results and Discussion
In this section, we numerically demonstrate the accuracy and efficiency of PIGen-SQD against SQD. To assess the performance of our approach we compute the ground state energies of three strongly correlated systems, (equilibrium O-H bond distance , bond angle ), () and () for different geometries using 6-31G basis set. With the frozen core approximation, , and consists of 24, 32 and 40 spin orbitals respectively which are mapped directly to 24, 32 and 40 qubits on IBM Heron R2 quantum processors as per the details discussed in Section IV. For Gibbs sampling process, the number of generated determinants is fixed throughout the iteration process at of the symmetry space dimension . The value is used in all the studies, unless mentioned otherwise.
@@ V.1 RBM Driven Sampling v/s Random Sampling
A crucial aspect of any such stochastic sampling driven subspace expansion method is the size of the search space explored by the ML model during the self-consistent configuration reconstruction process. Owing to the inherent stochasticity in the generative step, each macro-iteration inevitably produces configurations that have negligible or no contribution to the target state. These are collected into a blacklist to prevent their re-selection in subsequent iterations. Because the target CI vector is typically sparse, this blacklist grows much faster than the dominant (core) configuration space. As a result, the generated samples increasingly contain a significant fraction of unimportant configurations, causing the overall search space, which consists of both dominant and blacklisted configurations, to expand proportionally with iteration count. All of these quantitative measures are shown in Fig.3 where evidently, the RBM based method provides almost two orders of magnitude more accurate energy estimation while operating within a much smaller diagonalization space compared to random sampling. The associated search space dimension, which is a direct measure of the explored Hilbert space, is also substantially low for RBM driven generation whereas the same for random sampling process grows much more steeply. For random sampling to provide accurate results, it requires sufficiently large search space () which is impractical from a scalability perspective. Ideally the generative model should be trained in a way such that the blacklist space dimension is as low as possible, maximizing the number of dominant configurations during Hilbert space exploration. This concentrated search and configuration selection only from the relevant portion of the symmetry-restricted Hilbert space demonstrates an effective learning and configuration generation capability of the RBM.
@@
To numerically illustrate the degree of anchoring effect of the physics-informed initialization provided to hardware samples, Fig. 4 shows a comparative study regarding different rank of perturbative measures having different initialization to the fermionic state reconstruction process for different geometries of and . This analysis incorporates both symmetry-preserved hardware samples and perturbatively generated initializations with determinants up to ranks . Approximate perturbative estimates at each rank were obtained by including a subset of symbolically generated algebraic terms (see Appendix A, B). The plot demonstrates that ML-based configuration recovery converges much faster with better energy accuracy when initialized with excited configurations up to rank 4 compared to only hardware, (up to) rank 2 and rank 3 initializations. This is observed for both equilibrium and stretched geometries where strong correlation effects are more prevalent. The inset bar plots show the number of configurations required in the search space to reach chemical accuracy. In equilibrium geometries, determinants taken up to rank 4 provide an excellent starting point and hence requires substantially reduced dimensional search space to reach chemical accuracy. This ensures an effective learning by the generative model of the distribution of the target state basis vectors. For stretched geometries also, similar trend persists, though the relative difference between the search spaces narrows down compared to equilibrium geometries (see the inset bar plots) as the presence of strong correlation requires a substantial number of even higher rank determinants and more search space. However, even in very strongly correlated regime where the low order perturbative estimates may miss out a few important configurations, the generative learning algorithm is eventually able to pick those up, albeit with a slower macro-cyclic convergence and somewhat larger search space. Moreover, PIGen-SQD initialized with raw symmetry-preserved hardware samples shows similar accuracy and search space dimensions compared to rank 2 and rank 3 initializations in most of the cases under consideration in Fig. 4. This suggests that one needs to strike the right balance between classical physics-informed initialization and the generative procedure to an accurate and efficacious energy estimation. Toward this, from here onward, we will use for all PIGen-SQD calculations. However, one may choose a higher value of for more strongly correlated systems to accelerate convergence.
@@ V.3 Potential Energy Profile and State-Fidelity:
The dissociation curve of via symmetric bond stretching in 6-31G basis set with frozen core approximation (8 electrons in 12 spatial orbitals) is shown in Fig. 5 for PIGen-SQD, SQD and other reference methods such as HF, MP2, CISD and CCSD for comparison. Panel (A) of Fig.5 shows the potential energy surface and the associated energy difference in with FCI is shown in panel (B) along y-axis with the ratio of bond distance to equilibrium distance plotted along x-axes. The comparative study of the diagonalization dimension for SQD and PIGen-SQD is shown in panel (C) along with the associated search space, spin-particle number conserved symmetry space and core space dimension. SQD is performed in this case with sample size 500 and 1000 in 5 batches at each step of configuration recovery loop while PIGen-SQD results are obtained with ( here) of the symmetry space samples generated in each macro-cycle. For SQD, the minimum energy and maximum diagonalization dimension among all batches are plotted. SQD with sample size 500 or SQD(500) shows chemically accurate results near equilibrium geometry while it struggles for stretched bonds. Since increasing sample size includes more configurations in the diagonalization subspace, SQD(1000) performs well for the entire dissociation curve. On the other hand, PIGen-SQD() and PIGen-SQD() both shows results well within the chemical accuracy. However, the main aspect that distinguishes the efficiency of the two methods is the associated diagonalization dimension. In this study, while SQD(1000) requires almost around 84% of the entire symmetry space for diagonalization to provide chemically accurate results for the entire potential energy surface, PIGen-SQD(2%) does so using only of the same. Moreover, our method is systematically improvable by increasing during generative process as demonstrated in the plot with the PIGen-SQD(6%) that shows substantially improved accuracy compared to SQD(1000) throughout the dissociation curve particularly in the strongly correlated regime. Another important metric of our method is the search space that RBM explores during the self-consistent configuration recovery process. From Fig.5 it is evident that the search space for PIGen-SQD is around the core space dimension for for the entire dissociation curve while it slightly increases for providing better accuracy. However, in all the cases, even the search space dimension is well below that of SQD diagonalization dimension, showing an effective prediction of the underlying distribution of the dominant states in the Hilbert space. While PIGen-SQD can be systematically improved by tuning , in our subsequent studies, we will use as it strikes an optimal balance between the accuracy and the explored Hilbert search.
@@
Bond stretching of the () molecule is a paradigmatic approach to demonstrate electronic strong correlation effects. Fig. 6 shows the dissociation curve of in 6-31G basis with frozen core approximation resulting in 10 electrons in 16 spatial orbitals for both SQD and PIGen-SQD. The SQD results are obtained using sample size 3000 and 5000 with fixed batch size 5. While both SQD(3000) and SQD(5000) shows chemically accurate results near equilibrium geometries, it fails in stretched geometries having errors of the order of and with around 6 and 10 million determinants in the diagonalization space respectively. Contrarily, PIGen-SQD(2%) shows drastically improved results maintaining around accuracy throughout the dissociation curve, even under strong correlation at stretched geometries. Remarkably, PIGen-SQD needs only about 1-1.5 million determinants in the diagonalization subspace, which shows almost an order-of-magnitude reduction in the diagonalization subspace while providing more than an order-of-magnitude energy accuracy. The associated search space is well under the core space dimension near equilibirum geometries, while it increases to of the order of core space dimension for stretched geometries. However, the search space is still well below the diagonalization dimension of SQD(5000).
@@
Fig. 7 illustrates an approximate estimation of fidelity of the reconstructed wavefunctions by comparing the CI coefficients obtained from the different methods against FCI coefficients. The absolute CI coefficients obtained from FCI, SQD and PIGen-SQD are sorted in an ascending order and plotted along the x-axis for the two geometries of ((A), (B)) and ((C), (D)). For any given interval on the x-axis, the corresponding y-value indicates the number of coefficients falling within that range. The corresponding energy difference from FCI is given in the inset text for all the cases. For both SQD and PIGen-SQD shows similar order of accuracy though the number of configurations predicted by SQD is substantially large. Fig.7 (A, B) histograms show PIGen-SQD distribution has a peak toward the larger CI coefficients () representing a dominant subset of the FCI coefficients that provides major support to target state and contribute more toward accurate energy estimation. Better state fidelity for PIGen-SQD is more conspicuous for stretched geometries of in Fig. 7 (C, D) as one can see the substantially higher peak of PIGen-SQD curve toward the region with larger CI coefficients than SQD which has a flatter distribution over the entire range of values. This is reflected in the corresponding energy errors (see the inset text) where PIGen-SQD shows more than one order of magnitude better energy accuracy (0.3 compared to 4 obtained by SQD) under strong correlation. These examples demonstrate that PIGen-SQD faithfully reconstructs accurate high-fidelity states as compared to SQD leading to significantly improved energy predictions with substantially reduced requirements of HPC resources. The corresponding density plot in Fig.9 shows the predicted coefficients are sharply peaked around larger values with maximum probability for PIGen-SQD, making it a unimodal distribution. On the contrary, SQD and FCI curves are bimodal with a higher peak near practically zero coefficients due to sparsity. This further confirms the concentrated sampling of the dominant sector of the Hilbert space enabled by PIGen-SQD.

### v2

@@ III.2 Generative Machine Learning Based Fermionic State Reconstruction
where, is the partition function. Since explicit calculation of the partition function becomes computationally intractable, RBMs are usually trained on a set of vectors distributed according to some unknown probability distribution using the contrastive divergence (CD)[61] method which approximates the gradient of the log likelihood to learn the underlying distribution. RBM alternatively updates hidden and visible layers using conditional activation probabilities:
where, is the sigmoid function. By maximizing the log likelihood of the data , RBM uses stochastic approximation to update its parameters which is equivalent to minimizing the Kullback-Leibler divergence between the data distribution and model distribution. After training, a new set of vectors are generated according to the learnt distribution using Gibbs sampling[62] technique. For a detailed description of the working principle of RBM, the readers are referred to Refs.[61, 63, 64, 65, 66].
@@
Initial State Preparation and Sampling from Quantum Hardware: State preparation is done on a quantum hardware with an ansatz that has a substantial overlap with the ground state. Usually a PQC corresponding to a unitary is prepared that acts on HF state such that the prepared state is . This is followed by repeated measurements performed with a specific number of shots to obtain samples from the hardware that produce a measurement outcome with a probability distribution where each bitstring is a computational basis state.
@@
Micro Cycle: Each macro-cycle contains a set of micro-cycle iterations which is used to train RBM with the prepared input vector distribution. Trained RBM is used to generate new configurations using Gibbs sampling method. Toward this, a set of spin-particle conserved configurations, whose number is fixed at % of , are randomly sampled from the symmetry space and given to the Gibbs sampler which generate same number of spin-particle symmetry conserved configurations.
@@ IV Computational Details
All the numerical results in this work are obtained by in-house python based codes. MBPT calculations are done using two-body integrals and MP2 amplitudes obtained from PySCF[67] driver in Qiskit[68]. Three strongly correlated molecular systems , and in 6-31G basis with frozen core approximation involving 24, 32 and 40 spin orbitals correlating 8, 10 and 10 electrons respectively are studied in this work. The FCI, CCSD, CI with singles and doubles (CISD), MP2 and HF calculations were performed using PySCF with the frozen core approximation. Since FCI calculations are not feasible for , heat-bath configuration interaction (HCI)[39] calculations were done using PySCF to get reference energy values for accuracy comparison. The entire workflow involves multiple independent computational components orchestrated together as discussed below:
Quantum state preparation and hardware sample generation: State preparation on a quantum hardware is done using the LUCJ ansatz [45] (Eq. (6)) with CCSD parameters used for jastrow decomposition. No further optimization of the LUCJ parameters is performed. For 24 qubits experiment 250000 shots were used while both 32 and 40 qubit experiments were performed using shots. The circuits were created using qiskit-addon-sqd, which uses the ffsim[69] package to generate the LUCJ circuits. For each circuit, we start with some generic linear layout and invoke the VF2PostLayout transpiler pass to find an isomorphic zig-zag layout (which is required for LUCJ circuits) that conforms to the lowest noise profile as obtained from the backend noise information. The LUCJ circuit is placed on this layout using the qiskit transpiler. We executed all the circuits on ibm_kingston, which is a 156 qubits IBM Heron R2 processor. The layouts used for the circuits corresponding to each of the molecules are shown in Fig. 1.
@@
To obtain the physics-informed support to ground state via perturbative measures, initially available molecular integrals (in molecular orbital basis) are filtered from the original sparse integral tensor using a threshold . All surviving integral values, together with their associated index combinations, are collected and arranged into an array to construct following the integral index pattern (). The computational cost of triples selection step is governed by the most expensive term which can be performed efficiently using the symbolic index operation technique discussed in Sec.III. For this efficient generation we combine all unique outer indices and having only one index common between them (see Appendix B) with the constraints and respectively, stemming from pruned and . Such an operation does not require explicit summations over contractible indices, and just one common index suffices the inclusion of the corresponding outer (non-contractible) indices. At this stage, it is important to note that the hierarchical process of configuration selection through perturbation theory automatically ensures that the lower order terms, double excitations () from the HF reference, are already pruned which implies . This provides the flexibility to avoid the prohibitive cost of explicit tensor contraction, saving huge number of operations to be performed. A comparison of the number of operations required for the pruned elements in the numerical worst case scenario for triples and quadruples selection are shown in Fig.2 against the formal asymptotic scaling of the term , CCSD and CCSDT. Fig. 2 shows the numerical limiting case of symbolic tensor manipulation with pruned elements for triples selection requires almost two orders-of-magnitude less number of operations compared to the asymptotic cost of explicit computations. The actual implementational cost for the same using the symbolic technique requires much less number of operations than this in all practical cases. A similar approach can be followed for constructing perturbative quadruples as well by breaking the associated contraction into intermediates and performing the index-based symbolic calculations. The algebraic terms calculated for selection of quadruply excited configurations are given in Appendix A. The corresponding numerical limiting case scales almost similar to CCSD as demonstrated in Fig.2.
@@
While this way of selection of important high-rank excited configurations through their lowest order perturbative measures is not theoretically exact from a conventional many-body viewpoint, we argue that an absolute accuracy (of choosing the important configurations) is not also warranted for; rather, this measure gives us a strong additional support to the hardware generated configuration samples. Dominant determinants that are missed out (due to noise or perturbatively) would eventually get picked up through generative ML. Also, in a similar spirit, any determinant which are spuriously picked up would get dropped (blacklisted) during the iterative fermionic state reconstruction as will be shown in numerical demonstrations.
@@
Generative ML training and fermionic state reconstruction: In this study, we have used RBM with CD algorithm for training. The dimensions of visible and hidden layer are kept the same in this study which is equal to the number of spin orbitals, the learning rate is fixed at 0.001. The training distribution provided to the RBM is constructed by inflating the current configuration space obtained in a given macro-iteration, with each configuration replicated in proportion to its associated CI coefficient (excluding the HF configuration, as discussed in Sec. III.2). Within each RBM micro-cycle, three training iterations are performed using the CD algorithm to update the model parameters. During these updates, CD incorporates an intrinsic Gibbs-sampling step, for which we employ 20 Gibbs iterations. With number of training iterations (or epochs), intrinsic Gibbs sampling steps, inputs, visible layers and hidden layers, the number of operations is bounded by [70] During generation, only one Gibbs sampling step is used that takes a fixed number of randomly selected configurations from the correct spin-particle sector of the Hilbert space and generates the same number of configurations using the learned parameters. The number of samples used for Gibbs sampling and subsequent generation at each macro-cycle is fixed a priori by configurations of the symmetry space where the number is specified for the numerical studies in the result section (Sec. V). Such a training and generation process avoids over-training and retain a low degree of stochasticity in the model for better generation of new configurations as discussed further in Sec.V. The generated determinants are screened against a blacklist containing configurations with a complexity while the associated diagonalization scales as with determinants in the diagonalization space.[49, 71] For construction of the blacklist, the associated cutoff threshold is taken as and convergence threshold for the self-consistent loop is .
@@ V Results and Discussion
In this section, we numerically demonstrate the accuracy and efficiency of PIGen-SQD against SQD. To assess the performance of our approach we compute the ground state energies of three strongly correlated systems, (equilibrium O-H bond distance , bond angle ), () and () for different geometries using 6-31G basis set. With the frozen core approximation, , and consists of 24, 32 and 40 spin orbitals respectively which are mapped directly to 24, 32 and 40 qubits on IBM Heron R2 quantum processors as per the details discussed in Section IV. For Gibbs sampling process, the number of generated determinants is fixed throughout the iteration process at of the symmetry space dimension . The value is used in all the studies, unless explicitly mentioned otherwise.
@@ V.1 RBM Driven Sampling v/s Random Sampling
A crucial aspect of any such stochastic sampling driven subspace expansion method is the size of the search space explored by the ML model during the self-consistent configuration reconstruction process. Owing to the inherent stochasticity in the generative step, each macro-iteration inevitably produces configurations that have negligible or no contribution to the target state. These are collected into a blacklist to prevent their re-selection in subsequent iterations. Because the target CI vector is typically sparse, this blacklist grows much faster than the dominant configuration space. As a result, the generated samples increasingly contain a significant fraction of unimportant configurations, causing the overall search space, which consists of both dominant and blacklisted configurations, to expand proportionally with iteration count. All of these quantitative measures are shown in Fig.3 where evidently, the RBM based method provides almost two orders of magnitude more accurate energy estimation while operating within a much smaller diagonalization space compared to random sampling. The associated search space dimension, which is a direct measure of the explored Hilbert space, is also substantially low for RBM driven generation whereas the same for random sampling process grows much more steeply. For random sampling to provide accurate results, it requires sufficiently large search space () which is impractical from a scalability perspective. Ideally the generative model should be trained in a way such that the blacklist space dimension is as low as possible, maximizing the number of dominant configurations during Hilbert space exploration. This concentrated search and configuration selection only from the relevant portion of the symmetry-restricted Hilbert space demonstrates an effective learning and configuration generation capability of the RBM.
@@
To numerically illustrate the degree of anchoring effect of the physics-informed initialization provided to hardware samples, Fig. 4 shows a comparative study regarding different rank of perturbative measures having different initialization to the fermionic state reconstruction process for different geometries of and . This analysis incorporates both symmetry-preserved hardware samples and perturbatively generated initializations with determinants up to ranks . Approximate perturbative estimates at each rank were obtained by including a subset of symbolically generated algebraic terms (see Appendix A, B). The plot demonstrates that ML-based configuration recovery converges much faster with better energy accuracy when initialized with excited configurations up to rank 4 compared to only hardware, (up to) rank 2 and rank 3 initializations. This is observed for both equilibrium and stretched geometries where strong correlation effects are more prevalent. The inset bar plots show the number of configurations required in the search space to reach chemical accuracy. In equilibrium geometries, determinants taken up to rank 4 provide an excellent starting point and hence requires substantially reduced dimensional search space to reach chemical accuracy. This ensures an effective learning by the generative model of the distribution of the target state basis vectors. For stretched geometries also, similar trend persists, though the relative difference between the search spaces narrows down compared to equilibrium geometries (see the inset bar plots) as the presence of strong correlation requires a substantial number of even higher rank determinants and more search space. However, even in very strongly correlated regime where the low order perturbative estimates may miss out a few important configurations, the generative learning algorithm is eventually able to pick those up, albeit with a slower macro-cyclic convergence and somewhat larger search space. Moreover, initializations with raw symmetry-preserved hardware samples show similar accuracy and search space dimensions compared to rank 2 and rank 3 initializations in most of the cases under consideration in Fig. 4. This suggests that one needs to strike the right balance between classical physics-informed initialization and the generative procedure to an accurate and efficacious energy estimation. Toward this, from here onward, we will use for all PIGen-SQD calculations. However, one may choose a higher value of (with some selected algebraic terms that scale favorably) for more strongly correlated systems to accelerate convergence which would be explored further in future works.
@@ V.3 Potential Energy Profile and State-Fidelity:
The dissociation curve of via symmetric bond stretching in 6-31G basis set with frozen core approximation (8 electrons in 12 spatial orbitals) is shown in Fig. 5 for PIGen-SQD, SQD and other reference methods such as HF, MP2, CISD and CCSD for comparison. Panel (A) of Fig.5 shows the potential energy surface and the associated energy difference in with FCI is shown in panel (B) along y-axis with the ratio of bond distance to equilibrium distance plotted along x-axes. The comparative study of the diagonalization dimension for SQD and PIGen-SQD is shown in panel (C) along with the associated search space, spin-particle number conserved symmetry space and core space dimension. SQD is performed in this case with sample size 500 and 1000 in 5 batches at each step of configuration recovery loop while PIGen-SQD results are obtained with ( here) of the symmetry space samples generated in each macro-cycle. For SQD, the minimum energy and maximum diagonalization dimension among all batches are plotted. SQD with sample size 500 or SQD(500) shows chemically accurate results near equilibrium geometry while it struggles for stretched bonds. Since increasing sample size includes more configurations in the diagonalization subspace, SQD(1000) performs well for the entire dissociation curve. On the other hand, PIGen-SQD() and PIGen-SQD() both shows results well within the chemical accuracy. However, the main aspect that distinguishes the efficiency of SQD and PIGen-SQD is the associated diagonalization dimension. In this study, while SQD(1000) requires almost around 84% of the entire symmetry space for diagonalization to provide chemically accurate results for the entire potential energy surface, PIGen-SQD(2%) does so using only of the same. Moreover, our method is systematically improvable by increasing during generative process as demonstrated in the plot with the PIGen-SQD(6%) that shows substantially improved accuracy compared to SQD(1000) throughout the dissociation curve particularly in the strongly correlated regime. Another important metric of our method is the search space that RBM explores during the self-consistent configuration recovery process. From Fig.5 it is evident that the search space for PIGen-SQD is around the core space dimension for for the entire dissociation curve while it slightly increases for providing better accuracy. However, in all the cases, even the search space dimension is well below that of SQD diagonalization dimension, showing an effective prediction of the underlying distribution of the dominant states in the Hilbert space. While PIGen-SQD can be systematically improved by tuning , in our subsequent studies, we will use as it strikes an optimal balance between the accuracy and the explored Hilbert space.
@@
Bond stretching of the () molecule is a paradigmatic approach to demonstrate electronic strong correlation effects. Fig. 6 shows the dissociation curve of for both SQD and PIGen-SQD along with other standard methods in 6-31G basis with frozen core approximation resulting in 10 electrons in 16 spatial orbitals. The SQD results are obtained using sample size 3000 and 5000 with fixed batch size 5. For some stretched geometries CCSD did not converge and those points are not shown in the figure. While both SQD(3000) and SQD(5000) shows chemically accurate results near equilibrium geometries, it fails in stretched geometries having errors of the order of and with around 6 and 10 million determinants in the diagonalization space respectively. Contrarily, PIGen-SQD(2%) shows drastically improved results maintaining around accuracy throughout the dissociation curve, even under strong correlation at stretched geometries. Remarkably, PIGen-SQD needs only about 1-1.5 million determinants in the diagonalization subspace, which shows almost an order-of-magnitude reduction in the diagonalization dimension while providing more than an order-of-magnitude energy accuracy. The associated search space is well under the core space dimension near equilibirum geometries, while it increases to of the order of core space dimension for stretched geometries. However, the search space is still well below the diagonalization dimension of SQD(5000).
@@
Fig. 7 illustrates an approximate estimation of fidelity of the reconstructed wavefunctions by comparing the CI coefficients obtained from the different methods against FCI coefficients. The absolute CI coefficients obtained from FCI, SQD and PIGen-SQD are sorted in an ascending order and plotted along the x-axis for the two geometries of ((A), (B)) and ((C), (D)). For any given interval on the x-axis, the corresponding y-value indicates the number of coefficients falling within that range. The corresponding energy difference from FCI is given in the inset text for all the cases. For both SQD and PIGen-SQD shows similar order of accuracy though the number of configurations predicted by SQD is substantially large. Fig.7 (A, B) histograms show PIGen-SQD distribution has a peak toward the larger CI coefficients () representing a dominant subset of the FCI coefficients that provides major support to target state and contribute more toward accurate energy estimation. Better state fidelity for PIGen-SQD is more conspicuous for stretched geometries of in Fig. 7 (C, D) as one can see the substantially higher peak of PIGen-SQD curve toward the region with larger CI coefficients than SQD which has a flatter distribution over the entire range of values. This is reflected in the corresponding energy errors (see the inset text) where PIGen-SQD shows more than one order of magnitude better energy accuracy (0.3 compared to 4 obtained by SQD) under strong correlation. These examples demonstrate that PIGen-SQD faithfully reconstructs accurate high-fidelity states as compared to SQD leading to significantly improved energy predictions with substantially reduced requirements of HPC resources. The corresponding density plot in Fig.9 (see Appendix C) shows the predicted coefficients are sharply peaked around larger values with maximum probability for PIGen-SQD, making it a unimodal distribution. On the contrary, SQD and FCI curves are bimodal with a higher peak near practically zero coefficients due to sparsity. This further confirms the concentrated sampling of the dominant sector of the Hilbert space enabled by PIGen-SQD.
@@
It is important to mention in this aspect that while in all the cases under consideration PIGen-SQD shows up to 90 reduction in the diagonalization subspace dimension along with orders-of-magnitude more accurate energy estimations, the performance could be systematically improved further with increased shot budget, optimized hyper parameters and more robust machine learning models such as transformers.[72]